. Air at a temperature of 20°C passes through a heat exchanger at a

velocity of 40 m/s where its temperature is raised to 820°C. It then enters a turbine with same

velocity of 40 m/s and expands till the temperature falls to 620°C. On leaving the turbine, the air

is taken at a velocity of 55 m/s to a nozzle where it expands until the temperature has fallen to

510°C. If the air flow rate is 2.5 kg/s, calculate :

(i) Rate of heat transfer to the air in the heat exchanger ;

(ii) The power output from the turbine assuming no heat loss ;

(iii) The velocity at exit from the nozzle, assuming no heat loss.

Take the enthalpy of air as h = cpt, where cp is the specific heat equal to 1.005 kJ/kg°C and

t th


1
Expert's answer
2021-08-30T17:54:51-0400

Solution;

Given;

Temperature of air,"t_1=20\u00b0c"

Velocity of air, "C_1=40m\/s"

Temperature of air after passing the heat exchanger,"t_2=820\u00b0c"

Velocity of air at entry to the turbine,"C_2=40m\/s"

Temperature of air leaving the turbine,"t_3=620\u00b0c"

Velocity of air at entry to nozzle,"C_3=55m\/s"

Temperature of air after expansion through the nozzle,"t_4=510\u00b0c"

Air flow rate, "\dot{m}=2.5kg/s"

(i) Heat exchanger;

To find the rate of heat transfer;

From Energy equation given by;

"m(h_1+\frac{C_1^2}{2}+z_1g)+Q_{1-2}=m(h_2+\frac{C_2^2}{2}+z_2g)+W_{1-2}"

In which ;

"z_1=z_2"

"C_1,C_2" "=0"

"W_{1-2}=0"

Therefore ,the equation reduces to;

"mh_1+Q_{1-2}=mh_2"

"Q_{1-2}=m(h_2-h_1)"

"Q_{1-2}=mc_p(t_2-t_1)"

"Q_{1-2}=2.5\u00d71.005\u00d7(820-20)"

"Q_{1-2}=2010kJ\/s"

Answer;

2010kJ/s

(ii) Turbine;

Power output from Turbine;

Energy equation for turbine will be ;

"m(h_2+\frac{C_2^2}{2})=m(h_3+\frac{C_3^2}{2})+W_{2-3}"

Since;

"Q_{2-3}=0"

"z_1=z_2=z_3"

"W_{2-3}=m[(h_2-h_3)+(\frac{C_2^2-C_3^2}{2})]"

"W_{2-3}=m[c_p(t_2-t_3)+\frac{C_2^2-C_3^2}{2}]"

"W_{2-3}=2.5[1.005(820-620)+(\frac{40^2-55^2}{2×1000})]"

"W_{2-3}=2.5[201+0.7125]"

"W_{2-3}=504.3kJ\/s"

Answer;

504.3kW

(iii) Nozzle

Velocity at exit from the nozzle;

Energy Equation for the nozzle will be ;

"h_3+\frac{C_3^2}{2}=h_4+\frac{C_4^2}{2}"

Since;

"z_1=z_2"

"W_{3-4}=0"

"Q_{2-3}=0"

"\frac{C_4^2}{2}=c_p(t_3-t_4)+\frac{C_3^2}{2}"

"\frac{C_4^2}{2}=1.005(620-510)+\frac{55^2}{2×1000}"

"C_4=473m\/s"

Answer;

473m/s